OpenStudy (anonymous):

If anyone could help it would be greatly appreciated the number of visitors to the us nation parks in time is given by the formula n=-0.08t^2+1.64t+50, where n represents millions of visitors and t is in years since 1998. in what year do we expect 30 million tourists in our parks. round up to the next calendar year.

OpenStudy (campbell_st):

well set n = 30 and then solve for t so \[30 = -0.08t^2 + 1.64t + 50\] or you should use the general quadratic formula to solve the problem.

OpenStudy (campbell_st):

you need to use \[t = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = -0.08, b = 1.64 and c = 20 only consider the positive answer, which will need to be added to 1998 to find the year.